Assume that in the year 2007 stock and net fixed assets amounted to ksh 22 m and 30 m respectively sales amounted to 50 m. Calculate the stock and fixed assets as a percentage of sales.
1
Expert's answer
2018-10-29T16:35:08-0400

%stock assets=22/50=44%

%fixed assets=30/50=60%